# Robust Max Statistics for High-Dimensional Inference

### v1

@@ 3 Theory
  1. (i).
    For all $v\in\mathbb{R}^{p}$, the inequality $\big{\|}\big{\langle}v,X_{1}-\mathbf{E}(X_{1})\big{\rangle}\big{\|}_{L^{4+\delta}}\,\leq\,C\big{\|}\big{\langle}v,X_{1}-\mathbf{E}(X_{1})\big{\rangle}\big{\|}_{L^{2}}$ holds.
  2. (ii).
    For all $j=1,\dots,p$, the random variable $X_{1j}/\sigma_{j}$ has a density $f_{j}$ with respect to Lebesgue measure on $\mathbb{R}$ such that $\|f_{j}\|_{L^{\infty}}\leq C$.
  3. (iii).
@@
Remarks. All of the conditions in Assumption 1 are invariant to shifting $X_{1}\mapsto X_{1}+v$ for fixed $v\in\mathbb{R}^{p}$, and scaling $X_{1}\mapsto cX_{1}$ for fixed $c\neq 0$. The following paragraphs provide several examples that address each of the conditions (i)-(iv). Also, it is straightforward to combine the examples to construct a wide assortment of data-generating distributions that satisfy all of the conditions in Assumption 1 simultaneously.
Examples of $L^{4+\delta}$-$L^{2}$ moment equivalence. In recent years, moment assumptions similar to condition (i) have been adopted in many analyses of robust statistical methods for high-dimensional data lugosi2019mean ; ke2019user ; mendelson2020robust ; roy2021empirical ; abdalla2022covariance . As shown in Proposition 1 below, condition (i) is compatible with the classes of elliptical and separable models (also known as independent component models), which are widely used in areas such as multivariate analysis, random matrix theory, and signal processing (kotz2019continuous, ; Bai:2010, ; Comon:2010, ).
@@
  • (1).
    The observation $X_{1}$ is drawn from an elliptical distribution such that $\|\eta_{1}\|_{L^{4+\delta}}\lesssim\sqrt{p}$ and $\limsup_{t\to 0^{+}}\frac{1}{t}\mathbf{P}(\eta_{1}^{2}/p\leq t)\lesssim 1$. (Note that the latter condition holds if $\eta_{1}^{2}/p$ has a density $h_{p}$ with respect to Lebesgue measure such that $h_{p}(0)\lesssim 1$.)
  • (2).
    The observation $X_{1}$ is drawn from a separable distribution with $\max_{1\leq j\leq p}\|\zeta_{1j}\|_{L^{4+\delta}}$ $\lesssim$ $1$, and each random variable $\zeta_{1j}$ has a density $g_{j}$ with respect to Lebesgue measure such that $\max_{1\leq j\leq p}\|g_{j}\|_{L^{\infty}}\lesssim 1$.
@@
Principal components analysis. In the context of principal components analysis, it is common to assume that the sorted eigenvalues $\lambda_{1}(\Sigma)\geq\cdots\geq\lambda_{p}(\Sigma)$ of $\Sigma$ satisfy $\lambda_{j}(\Sigma)\lesssim j^{-\gamma}$ for some $\gamma>0$, and it can be shown that this implies $\sigma_{(j)}^{2}\lesssim j^{-2\beta}$ for some other decay parameter $\beta>0$ (lopes2020bootstrapping, , Proposition 2.1).
@@
Mean-variance proportionality. Another scenario where variance decay arises is when the coordinate-wise means and variances are connected by a proportionality relationship of the form $\sigma_{j}^{2}\propto|\mu_{j}|^{\gamma}$, for some fixed exponent $\gamma>0$. This occurs within many sub-families of distributions, including Gamma, Weibull, inverse Gaussian, lognormal and Pareto. In applications that involve sparse modelling of high-dimensional mean vectors, a classical assumption is that the sorted coordinate-wise means have a decay profile (Johnstone, ), and thus, when such a proportionality relationship with holds, it follows that variance decay must also occur.
@@
Functional data analysis. One more set of examples is related to functional data analysis, where function-valued observations $\Psi_{1},\dots,\Psi_{n}$ in a Hilbert space are often studied through their projections under a finite number of orthonormal basis functions $\{\phi_{j}\}_{1\leq j\leq p}$. That is, the $i$th projected observation has the form $X_{i}=(\langle\Psi_{i},\phi_{1}\rangle,\dots,\langle\Psi_{i},\phi_{p}\rangle)\in\mathbb{R}^{p}$. In connection with our work, the important point is that under standard assumptions in functional data analysis, it can be shown that the sorted coordinate-wise variances of $X_{i}$ have a decay profile (lopes2020bootstrapping, ). In fact, this occurs even when the random functions $\Psi_{1},\dots,\Psi_{n}$ have rough sample paths, which we illustrate empirically in Figure 2.
@@
Examples of correlation matrices. To interpret the condition (iv), it should be noted that the inequality $\|R(l_{n})\|_{F}^{2}\leq l_{n}^{2}$ always holds, since $\|A\|_{F}^{2}\leq\textup{tr}(A)^{2}$ holds for any positive semidefinite matrix $A$. So, in this sense, condition (iv) is quite mild, as $C$ may be taken to be arbitrarily large. Moreover, the correlation structure of the variables indexed by $\{1,\dots,p\}\setminus J(l_{n})$ is completely unrestricted. With regard to the constant 24 appearing in the definition of $l_{n}$, it has no special importance, and is used for theoretical convenience. Below, we describe several classes of $p\times p$ correlation matrices $R=R(p)$ for which the sub-matrix $R(l_{n})$ satisfies condition (iv).
@@
Decaying correlation functions. Let $\rho:[0,\infty)\rightarrow[0,1]$ be any continuous convex function satisfying $\rho(0)=1$, and $\rho(t)\leq ct^{-\gamma}$ for some fixed constants $c>0,\gamma>0$, and all $t\geq 0$. By Pólya’s criterion (polya1949remarks, ), a matrix whose $ij$ entry is defined by $\rho(|i-j|)$ is a correlation matrix that satisfies condition (iv). Correlation matrices of this type include many well-known examples, such as those of the autoregressive and banded types, e.g.  $R_{ij}=r^{|i-j|}$ for some fixed $r\in(0,1)$, and $R_{ij}=\max\big{\{}0,1-\frac{|i-j|}{b}\big{\}}$ for some fixed $b>0$.
@@
Diverging operator norm. If the operator norm of $R$ satisfies $\|R\|_{\textup{op}}\leq Cl_{n}^{\frac{1}{2}-\frac{1}{2C}}$, then Assumption 1(iv) holds. This can be seen by noting that $\|R(l_{n})\|_{F}^{2}\leq l_{n}\|R(l_{n})\|_{\textup{op}}^{2}\leq l_{n}\|R\|_{\textup{op}}^{2}$. In particular, since $l_{n}$ increases with $n$ and $p$, this shows that condition (iv) can hold even when the operator norm of $R$ diverges asymptotically.
@@
Block structure. Suppose that $R$ is formed by concatenating $k$ correlation matrices along its diagonal, with sizes $\nu_{1}\times\nu_{1},\dots,\nu_{k}\times\nu_{k}$, so that $\nu_{1}+\cdots+\nu_{k}=p$. If the condition $\max\{\nu_{1},\dots,\nu_{k}\}\leq Cl_{n}^{1-\frac{1}{C}}$ holds, then so does condition (iv). This follows from the observation that no row of $R$ can have a squared $\ell_{2}$ norm larger than $\max\{\nu_{1},\dots,\nu_{k}\}$, and so $\|R(l_{n})\|_{F}^{2}\leq l_{n}\max\{\nu_{1},\dots,\nu_{k}\}$.
@@
Convex combinations and permutations. If $R$ and $R^{\prime}$ denote any correlation matrices corresponding to the previous examples, then for any $t\in[0,1]$, the correlation matrix $tR+(1-t)R^{\prime}$ satisfies condition (iv). Furthermore, if $\Pi$ is a $p\times p$ permutation matrix, then $\Pi R\Pi^{\top}$ is also a correlation matrix that satisfies condition (iv). These operations considerably extend the examples that have decaying correlation functions or block structure.
@@
$\small\sup_{s\in\mathbb{R}}\Big{|}\mathbf{P}(\mathcal{M}_{n}\leq s)-\mathbf{P}(\mathcal{M}_{n}^{*}\leq s|X)\Big{|}\ \leq\ c\ \!n^{-\frac{1}{2}+\epsilon}$
@@ 4 Numerical experiments
This section addresses the practical performance of the proposed method in the contexts of Euclidean and functional data with heavy tails. In addition, we include an application to heavy-tailed financial data, as well as performance comparisons with the method for robust inference proposed in (Fan:2023, ).
@@ 4.1 Euclidean data
Design of experiments. For each pair $(n+m_{n},p)$ in the set $\{500\}\times\{100,500,1000\}$, and each of the data-generating distributions described below, we performed 500 Monte Carlo trials. Due to the fact that the proposed method and the method in (Fan:2023, ) are shift invariant, the mean vector $(\mu_{1},\dots,\mu_{p})$ was always chosen to be the zero vector without loss of generality. In all trials, the proposed confidence intervals ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ were constructed according to (7), using 500 bootstrap samples, and using choices of 90% and 95% for the nominal simultaneous coverage probability $1-\alpha$. Also, the proposed method was always applied with truncation parameters set to $\hat{t}_{j}=\sqrt{n}\hat{\sigma}_{j}$ for all $j=1,\dots,p$, and with the partial standardization parameter set to $\tau=0.9$. Lastly, in all trials, the number of hold-out observations was set to $m_{n}=50$, and the block length for the mom estimates was set to $\ell_{n}=10$.
@@
Elliptical distribution. The elliptical observations have the form $X_{1}=\mu+\eta_{1}\Sigma^{1/2}U_{1}$, where $U_{1}\in\mathbb{R}^{p}$ is uniformly distributed on the unit sphere, and $\eta_{1}\geq 0$ is a random variable that is independent of $U_{1}$ such that $3\eta_{1}^{2}/(2p)$ follows an $F$ distribution with $p$ and 6 degrees of freedom. This distribution for $X_{1}$ is more commonly known as a multivariate $t$-distribution on 6 degrees of freedom (muirhead2009aspects, ).
@@
Discussion of results. In Table 1, we report empirical simultaneous coverage probabilities and width measures, for both the proposed method (denoted PM) and the method based on the Hodges-Lehmann estimator (denoted HL) developed in (Fan:2023, ). The simultaneous coverage probabilities were computed as the fraction of the 500 Monte Carlo trials in which all $p$ intervals of a given method contained the corresponding parameters $\mu_{1},\dots,\mu_{p}$. The width measure was computed as the median width of the $p$ intervals, averaged over the 500 trials, and it is shown in parentheses below the simultaneous coverage probabilities.
@@
With regard to the width measure, Table 1 shows that the intervals produced by PM are much tighter than those produced by HL, sometimes by a factor of 4 or more. This is to be expected, because the HL intervals have equal widths across all $p$ coordinates, whereas PM adapts to the scale of each covariate and takes advantage of variance decay. This difference between the methods is also reflected in another pattern—which is that the width measure for PM decreases as $p$ becomes increases, whereas the width measure for HL stays essentially constant as $p$ increases, since HL uses unstandardized covariates.
Table 1: Comparison of simultaneous coverage probability and confidence interval width: PM refers to the proposed method, and HL refers to the method based on the Hodges-Lehmann estimator proposed in (Fan:2023, ).
@@ 4.2 Functional data
Beyond heavy-tailed Euclidean data, our approach to inference with robust max statistics can be applied to heavy-tailed functional data. Below, we study the problem of detecting a non-zero drift in functional observations that arise from geometric Brownian motion (GBM). This is a heavy-tailed stochastic process in the sense that its marginals are lognormal, which is a standard example of heavy-tailed univariate distribution (Foss, ; nair, ). The sample paths of GBM present additional challenges from the standpoint of functional data analysis, because they are rough. Furthermore, GBM is of broad interest in financial applications, where it is widely used for modelling securities prices (Oksendal, ).
@@
where $\mu(t)$ is fixed real-valued function on $[0,1]$ such that $S(t)$ resides in $L^{2}[0,1]$ almost surely, and $h\geq 0$ is a fixed parameter that measures the “distance” from the null hypothesis of zero drift that occurs when $h=0$. The parameters $\mu(t)$ and $\varsigma_{0}$ are treated as unknown. Under these conditions, we are interested in using a dataset $S_{1}(t),\dots,S_{n+m_{n}}(t)$ of i.i.d. samples of $S(t)$ to address the hypothesis testing problem
$\small\mathsf{H}_{0}:h=0\text{ \ \ \ vs. \ \ \ }\mathsf{H}_{1}:h>0.$ (9)
@@
$\small\mathsf{H}_{0,j}:\beta_{j}=0\text{ \ \ \ for \ \ \ }j=1,\dots,p,$ (10)
@@
If we define $X_{i}\in\mathbb{R}^{p}$ to contain the first $p$ coefficients of $S_{i}(t)-1$ with respect to $\{\phi_{j}(t)\}_{j\geq 1}$, then it follows that $\mathbf{E}(X_{i})=(\beta_{1},\dots,\beta_{p})$ for every $i=1,\dots,n+m_{n}$. Hence, we may apply our proposed method from Section 2 to $X_{1},\dots,X_{n+m_{n}}$ in order to construct confidence intervals ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\beta_{1},\dots,\beta_{p}$ with a nominal simultaneous coverage probability of $1-\alpha$. Altogether, this means that if we reject $\mathsf{H}_{0}$ when any of these intervals exclude 0, then this rejection rule corresponds to a test with a nominal level of at most $\alpha$.
@@
Design of experiments. To construct four natural choices of the pair $(\mu(t),\varsigma_{0})$, we used historical data for the stocks of Apple, Nvidia, Moderna, and JPMorgan. With regard to the choice of $\varsigma_{0}^{2}$, note that if a stock price is modeled as a realization of $S(t)$, then the pointwise variance of the cumulative log return is $\textup{var}((h\mu(t)-\varsigma_{0}^{2}/2)t+\varsigma_{0}W(t))=\varsigma_{0}^{2}t$, and the time average of this quantity over the unit interval is $\varsigma_{0}^{2}/2$. (See (ruppertstatistics, ) for additional background.) Accordingly, for each of the four stocks, we selected $\varsigma_{0}$ such that $\varsigma_{0}^{2}/2=\int_{0}^{1}s^{2}(t)\,\mathrm{d}t$, where $s^{2}(t)$ is the sample pointwise variance of the cumulative log return of the stock over many disjoint periods of unit length. (The data were prepared in the same manner as described in Section 4.3 below, except that a period of unit length here was taken to be 100 minutes.) Next, the selection of $\mu(t)$ was based on the fact that if a stock price is modeled with $S(t)$, then the pointwise expected cumulative log return is $(h\mu(t)-\varsigma_{0}^{2}/2)t$. Therefore, for a given value of $h$ and $\varsigma_{0}$, we defined $\mu(t)$ to satisfy $(h\mu(t)-\varsigma_{0}^{2}/2)t=\bar{R}(t)$, where $\bar{R}(t)$ is the sample average of the cumulative log return curves of a given stock over many disjoint periods of unit length.
@@
For each choice of the four choices of $(\mu(t),\varsigma_{0})$, and each value of $h$ in an equispaced grid, the following procedure was repeated in 500 Monte Carlo trials. We generated i.i.d. realizations $S_{1}(t),\dots,S_{n+m_{n}}(t)$ of the sample path defined in (8) with $n+m_{n}=300$ and $m_{n}=30$. For each of these sets of functional observations, we constructed simultaneous $(1-\alpha)$-confidence intervals ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for the parameters $\beta_{1},\dots,\beta_{p}$, as described above, with $p=100$ and $\alpha=5\%$. Whenever any of these intervals excluded 0, a rejection was recorded, and the rejection rate among the 500 trials was plotted as a function of $h$ in Figures 3(a)-3(d). The corresponding rejection rate based on the simultaneous confidence intervals proposed in the paper (Fan:2023, ) was also plotted in the same way. In all four figures, the nominal level of $\alpha=5\%$ is marked with a dashed horizontal line.
@@
To illustrate the characteristics of the simulated functional data, ten realizations of $S(t)$ based on $h=0$ with $\varsigma_{0})$ corresponding to Apple stock are plotted in Figure 2. In the same setting, Figure 2 displays estimates of the sorted values $\sigma_{(1)}\geq\cdots\geq\sigma_{(p)}$, where $\sigma_{j}^{2}$ is the variance of the $j$th Fourier coefficient of $S_{1}(t)$. In particular, Figure 2 shows that that the theoretical variance decay profile considered in Section 3 agrees well with the empirical behavior of GBM.

### v2

@@ 3 Theory
  1. (i)
    For all $v\in\mathbb{R}^{p}$, $\big\|\big\langle v,X_{1}-\mathbf{E}(X_{1})\big\rangle\big\|_{L^{4+\delta}}\,\leq\,C\big\|\big\langle v,X_{1}-\mathbf{E}(X_{1})\big\rangle\big\|_{L^{2}}$ holds.
  2. (ii)
    For all $j=1,\dots,p$, the random variable $X_{1j}/\sigma_{j}$ has a Lebesgue density $f_{j}$ such that $\|f_{j}\|_{L^{\infty}}\leq C$.
  3. (iii)
@@
Remarks. All of the conditions in Assumption 1 are invariant to shifting $X_{1}\mapsto X_{1}+v$ for fixed $v\in\mathbb{R}^{p}$, and scaling $X_{1}\mapsto cX_{1}$ for fixed $c\neq 0$. The following paragraphs provide several examples that address each of the conditions ((i))-((iv)). Also, it is straightforward to combine the examples to construct a wide assortment of data-generating distributions that satisfy all of the conditions in Assumption 1 simultaneously.
Examples of $L^{4+\delta}$-$L^{2}$ moment equivalence. In recent years, moment assumptions similar to condition ((i)) have been adopted in many analyses of robust statistical methods for high-dimensional data lugosi2019mean; ke2019user; mendelson2020robust; roy2021empirical; abdalla2022covariance. As shown in Proposition 1 below, condition ((i)) is compatible with the classes of elliptical and separable models (also known as independent component models), which are widely used in areas such as multivariate analysis, random matrix theory, and signal processing (kotz2019continuous; Bai:2010; Comon:2010).
@@
  • (1)
    The observation $X_{1}$ is drawn from an elliptical distribution such that $\|\eta_{1}\|_{L^{4+\delta}}\lesssim\sqrt{p}$, and the random variable $X_{11}/\sigma_{1}$ has a Lebesgue density $f_{1}$ such that $\|f_{1}\|_{L^{\infty}}\lesssim 1$.
  • (2)
    The observation $X_{1}$ is drawn from a separable distribution with
    $\max_{1\leq j\leq p}\|\zeta_{1j}\|_{L^{4+\delta}}$ $\lesssim$ 11, and each random variable $\zeta_{1j}$ has a Lebesgue density $g_{j}$ such that $\max_{1\leq j\leq p}\|g_{j}\|_{L^{\infty}}\lesssim 1$.
@@
Principal components analysis. In the context of principal components analysis, it is common to assume that the sorted eigenvalues $\lambda_{1}(\Sigma)\geq\cdots\geq\lambda_{p}(\Sigma)$ of $\Sigma$ satisfy $\lambda_{j}(\Sigma)\lesssim j^{-\gamma}$ for some $\gamma>0$, and it can be shown that this implies $\sigma_{(j)}^{2}\lesssim j^{-2\beta}$ for some other decay parameter $\beta>0$ (lopes2020bootstrapping, Proposition 2.1).
@@
Mean-variance proportionality. Another scenario where variance decay arises is when the coordinate-wise means and variances are connected by a proportionality relationship of the form $\sigma_{j}^{2}\propto|\mu_{j}|^{\gamma}$, for some fixed exponent $\gamma>0$. This occurs within many sub-families of distributions, including Gamma, Weibull, inverse Gaussian, and Pareto. In applications that involve sparse modelling of high-dimensional mean vectors, a classical assumption is that the sorted coordinate-wise means have a decay profile (Johnstone), and thus, when such a proportionality relationship holds, it follows that variance decay must also occur.
@@
Functional data analysis. One more set of examples is related to functional data analysis, where function-valued observations $\Psi_{1},\dots,\Psi_{n}$ in a Hilbert space are often studied through their projections under a finite number of orthonormal basis functions $\{\phi_{j}\}_{1\leq j\leq p}$. That is, the $i$th projected observation has the form $X_{i}=(\langle\Psi_{i},\phi_{1}\rangle,\dots,\langle\Psi_{i},\phi_{p}\rangle)\in\mathbb{R}^{p}$. In connection with our work, the important point is that under standard assumptions in functional data analysis, it can be shown that the sorted coordinate-wise variances of $X_{i}$ have a decay profile (lopes2020bootstrapping). In fact, this occurs even when the random functions $\Psi_{1},\dots,\Psi_{n}$ have rough sample paths, which we illustrate empirically in Figure 2.
@@
Examples of correlation matrices. To interpret the condition ((iv)), it should be noted that the inequality $\|R(l_{n})\|_{F}^{2}\leq l_{n}^{2}$ always holds, since $\|A\|_{F}^{2}\leq\textup{tr}(A)^{2}$ holds for any positive semidefinite matrix $A$. So, in this sense, condition ((iv)) is quite mild, as $C$ may be taken to be arbitrarily large. Moreover, the correlation structure of the variables indexed by $\{1,\dots,p\}\setminus J(l_{n})$ is completely unrestricted. With regard to the constant 24 appearing in the definition of $l_{n}$, it has no special importance, and is used for theoretical convenience. Below, we describe several classes of $p\times p$ correlation matrices $R=R(p)$ for which the sub-matrix $R(l_{n})$ satisfies condition ((iv)).
@@
Decaying correlation functions. Let $\rho\mathrel{\mathop{\ordinarycolon}}[0,\infty)\rightarrow[0,1]$ be any continuous convex function satisfying $\rho(0)=1$, and $\rho(t)\leq ct^{-\gamma}$ for some fixed constants $c>0,\gamma>0$, and all $t\geq 0$. By Pólya’s criterion (polya1949remarks), a matrix whose $ij$ entry is defined by $\rho(|i-j|)$ is a correlation matrix that satisfies condition ((iv)). Correlation matrices of this type include many well-known examples, such as those of the autoregressive and banded types, e.g.,  $R_{ij}=r^{|i-j|}$ for some fixed $r\in(0,1)$, and $R_{ij}=\max\big\{0,1-\frac{|i-j|}{b}\big\}$ for some fixed $b>0$.
@@
Diverging operator norm. If the operator norm of $R$ satisfies $\|R\|_{\textup{op}}\leq Cl_{n}^{\frac{1}{2}-\frac{1}{2C}}$, then Assumption 1((iv)) holds. This can be seen by noting that $\|R(l_{n})\|_{F}^{2}\leq l_{n}\|R(l_{n})\|_{\textup{op}}^{2}\leq l_{n}\|R\|_{\textup{op}}^{2}$. In particular, since $l_{n}$ increases with $n$ and $p$, this shows that condition ((iv)) can hold even when the operator norm of $R$ diverges asymptotically.
@@
Block structure. Suppose that $R$ is formed by concatenating $k$ correlation matrices along its diagonal, with sizes $\nu_{1}\times\nu_{1},\dots,\nu_{k}\times\nu_{k}$, so that $\nu_{1}+\cdots+\nu_{k}=p$. If the condition $\max\{\nu_{1},\dots,\nu_{k}\}\leq Cl_{n}^{1-\frac{1}{C}}$ holds, then so does condition ((iv)). This follows from the observation that no row of $R$ can have a squared $\ell_{2}$ norm larger than $\max\{\nu_{1},\dots,\nu_{k}\}$, and so $\|R(l_{n})\|_{F}^{2}\leq l_{n}\max\{\nu_{1},\dots,\nu_{k}\}$.
@@
Convex combinations and permutations. If $R$ and $R^{\prime}$ denote any correlation matrices corresponding to the previous examples, then for any $t\in[0,1]$, the correlation matrix $tR+(1-t)R^{\prime}$ satisfies condition ((iv)). Furthermore, if $\Pi$ is a $p\times p$ permutation matrix, then $\Pi R\Pi^{\top}$ is also a correlation matrix that satisfies condition ((iv)). These operations considerably extend the examples that have decaying correlation functions or block structure.
@@
$\sup_{s\in\mathbb{R}}\Big|\mathbf{P}(\mathcal{M}_{n}\leq s)-\mathbf{P}(\mathcal{M}_{n}^{*}\leq s|X)\Big|\ \leq\ c\ \!n^{-\frac{1}{2}+\epsilon}$
@@ 4 Numerical experiments
This section addresses the practical performance of the proposed method in the contexts of Euclidean and functional data with heavy tails. In addition, we include the performance comparisons with the method for robust inference proposed in (Fan:2023).
@@ 4.1 Euclidean data
Design of experiments. For each pair $(n+m_{n},p)$ in the set $\{500\}\times\{100,500,1000\}$, and each of the data-generating distributions described below, we performed 500 Monte Carlo trials. Due to the fact that the proposed method and the method in (Fan:2023) are shift invariant, the mean vector $(\mu_{1},\dots,\mu_{p})$ was always chosen to be the zero vector without loss of generality. In all trials, the proposed confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ were constructed according to (7), using 500 bootstrap samples, and using choices of 90% and 95% for the nominal simultaneous coverage probability $1-\alpha$. Also, the proposed method was always applied with truncation parameters set to $\hat{t}_{j}=\sqrt{n}\hat{\sigma}_{j}$ for all $j=1,\dots,p$, and with the partial standardization parameter set to $\tau=0.9$. Lastly, in all trials, the number of hold-out observations was set to $m_{n}=50$, and the block length for the mom estimates was set to $\ell_{n}=10$.
@@
Elliptical distribution. The elliptical observations have the form $X_{1}=\mu+\eta_{1}\Sigma^{1/2}U_{1}$, where $U_{1}\in\mathbb{R}^{p}$ is uniformly distributed on the unit sphere, and $\eta_{1}\geq 0$ is a random variable that is independent of $U_{1}$ such that $3\eta_{1}^{2}/(2p)$ follows an $F$ distribution with $p$ and 6 degrees of freedom. This distribution for $X_{1}$ is more commonly known as a multivariate $t$-distribution on 6 degrees of freedom (muirhead2009aspects).
@@
Discussion of results. In Table 1, we report empirical simultaneous coverage probabilities and width measures, for both the proposed method (denoted PM) and the method based on the Hodges-Lehmann estimator (denoted HL) developed in (Fan:2023). The simultaneous coverage probabilities were computed as the fraction of the 500 Monte Carlo trials in which all $p$ intervals of a given method contained the corresponding parameters $\mu_{1},\dots,\mu_{p}$. The width measure was computed as the median width of the $p$ intervals, averaged over the 500 trials, and it is shown in parentheses below the simultaneous coverage probabilities.
@@
With regard to the width measure, Table 1 shows that the intervals produced by PM are much tighter than those produced by HL, sometimes by a factor of 4 or more. This is to be expected, because the HL intervals have equal widths across all $p$ coordinates, whereas PM adapts to the scale of each covariate and takes advantage of variance decay. This difference between the methods is also reflected in another pattern—which is that the width measure for PM decreases as $p$ increases, whereas the width measure for HL stays essentially constant as $p$ increases, since HL uses unstandardized covariates.
Table 1: Comparison of simultaneous coverage probability and confidence interval width: PM refers to the proposed method, and HL refers to the method based on the Hodges-Lehmann estimator proposed in (Fan:2023).
@@ 4.2 Functional data
Beyond heavy-tailed Euclidean data, our approach to inference with robust max statistics can be applied to heavy-tailed functional data. Below, we study the problem of detecting a non-zero drift in functional observations that arise from geometric Brownian motion (GBM). This is a heavy-tailed stochastic process in the sense that its marginals are lognormal, which is a standard example of heavy-tailed univariate distribution (Foss; nair). The sample paths of GBM present additional challenges from the standpoint of functional data analysis, because they are rough. Furthermore, GBM is of broad interest in financial applications, where it is widely used for modelling securities prices (Oksendal).
@@
where $\mu(t)$ is a fixed real-valued function on $[0,1]$ such that $S(t)$ resides in $L^{2}[0,1]$ almost surely, and $h\geq 0$ is a fixed parameter that measures the “distance” from the null hypothesis of zero drift that occurs when $h=0$. The parameters $\mu(t)$ and $\varsigma_{0}$ are treated as unknown. Under these conditions, we are interested in using a dataset $S_{1}(t),\dots,S_{n+m_{n}}(t)$ of i.i.d. samples of $S(t)$ to address the hypothesis testing problem
$\small\mathsf{H}_{0}\mathrel{\mathop{\ordinarycolon}}h=0\text{ \ \ \ vs. \ \ \ }\mathsf{H}_{1}\mathrel{\mathop{\ordinarycolon}}h>0.$ (9)
@@
$\small\mathsf{H}_{0,j}\mathrel{\mathop{\ordinarycolon}}\beta_{j}=0\text{ \ \ \ for \ \ \ }j=1,\dots,p,$ (10)
@@
If we define $X_{i}\in\mathbb{R}^{p}$ to contain the first $p$ coefficients of $S_{i}(t)-1$ with respect to $\{\phi_{j}(t)\}_{j\geq 1}$, then it follows that $\mathbf{E}(X_{i})=(\beta_{1},\dots,\beta_{p})$ for every $i=1,\dots,n+m_{n}$. Hence, we may apply our proposed method from Section 2 to $X_{1},\dots,X_{n+m_{n}}$ in order to construct confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\beta_{1},\dots,\beta_{p}$ with a nominal simultaneous coverage probability of $1-\alpha$. Altogether, this means that if we reject $\mathsf{H}_{0}$ when any of these intervals exclude 0, then this rejection rule corresponds to a test with a nominal level of at most $\alpha$.
@@
Design of experiments. To construct four natural choices of the pair $(\mu(t),\varsigma_{0})$, we used historical data for the stocks of Apple, Nvidia, Moderna, and JPMorgan, to be described later in the paragraph labeled ‘preparation of stock data’. With regard to the choice of $\varsigma_{0}^{2}$, note that if a stock price is modeled as a realization of $S(t)$, then the pointwise variance of the cumulative log return is $\textup{var}((h\mu(t)-\varsigma_{0}^{2}/2)t+\varsigma_{0}W(t))=\varsigma_{0}^{2}t$, and the time average of this quantity over the unit interval is $\varsigma_{0}^{2}/2$. (See (ruppertstatistics) for additional background.) Accordingly, for each of the four stocks, we selected $\varsigma_{0}$ such that $\varsigma_{0}^{2}/2=\int_{0}^{1}s^{2}(t)\,\mathrm{d}t$, where $s^{2}(t)$ is the sample pointwise variance of the cumulative log return of the stock over many disjoint periods of unit length. Next, the selection of $\mu(t)$ was motivated by the fact that if a stock price is modeled with $S(t)$, and if $h=1$, then the pointwise expected cumulative log return is $(\mu(t)-\varsigma_{0}^{2}/2)t$. For a given value of $\varsigma_{0}$, we defined $\mu(t)$ to satisfy $(\mu(t)-\varsigma_{0}^{2}/2)t=\bar{R}(t)$, where $\bar{R}(t)$ is the sample average of the cumulative log return curves of a given stock over many disjoint periods of unit length.
@@
For each of the four choices of $(\mu(t),\varsigma_{0})$, and each value of $h$ in an equispaced grid, the following procedure was repeated in 500 Monte Carlo trials. We generated i.i.d. realizations $S_{1}(t),\dots,S_{n+m_{n}}(t)$ of the sample path defined in (8) with $n+m_{n}=300$ and $m_{n}=30$. For each of these sets of functional observations, we constructed simultaneous $(1-\alpha)$-confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for the parameters $\beta_{1},\dots,\beta_{p}$, as described above, with $p=100$ and $\alpha=5\%$. Whenever any of these intervals excluded 0, a rejection was recorded, and the rejection rate among the 500 trials was plotted as a function of $h$ in Figures 3(a)-3(d). The corresponding rejection rate based on the simultaneous confidence intervals proposed in the paper (Fan:2023) was also plotted in the same way. In all four figures, the nominal level of $\alpha=5\%$ is marked with a dashed horizontal line.
@@
To illustrate the characteristics of the simulated functional data, ten realizations of $S(t)$ based on $h=0$ with $\varsigma_{0}$ corresponding to Apple stock are plotted in Figure 2. In the same setting, Figure 2 displays estimates of the sorted values $\sigma_{(1)}\geq\cdots\geq\sigma_{(p)}$, where $\sigma_{j}^{2}$ is the variance of the $j$th Fourier coefficient of $S_{1}(t)$. In particular, Figure 2 shows a clear variance decay profile.